# Polyvalent Machine-Learned Potential for Cobalt: from Bulk to Nanoparticles

### v1

@@ Abstract
We present the development of a quadratic Spectral Neighbor Analysis Potential (q-SNAP) for ferromagnetic cobalt and its applications to bulk phases, surfaces, and nanoparticles. Trained on Density Functional Theory calculations using the Perdew-Burke-Ernzerhof (DFT-PBE) functional, this machine-learned potential enables simulations of large systems over extended time scales across a wide range of temperatures and pressures at near DFT accuracy. It is validated by closely reproducing the phonon dispersions of hexagonal close-packed (hcp) and face-centered cubic (fcc) Co, surface energies, and the relative stability of nanoparticles of various shapes. Thermal expansion and the melting point of Co computed with this potential are close to experimental values. Furthermore, this machine-learned potential goes beyond the capabilities of simpler N-body potentials by capturing nuanced properties such as vacancy formation energies on nanoparticle vertices. This accuracy and versatility make the potential suitable for a wide range of applications, including catalysis.
@@ I Introduction
Each of these new interatomic potentials is founded on a descriptor of atomic environments and a model that links them to local energies. Descriptors are mathematical entities invariant under translation, rotation, and permutation of atoms of the same type. A diverse array of descriptors exists, such as the bispectrum (BSO(4)) [17], the Smooth Overlap of Atomic Positions (SOAP) [18], and the Atomic Cluster Expansion (ACE) [19]. Likewise, models span from simple linear regression to multi-layer neural networks [17, 20, 21, 19, 22]. Thompson et al. [20] demonstrated a linear relationship between bispectrum descriptors and system observables (energy, forces and stress), leading to the introduction of the Spectral Neighbor Analysis Potential (SNAP). Later, they showed that adding quadratic terms to the model (leading to a q-SNAP) offers improved accuracy [23].
In this article, we present the construction of a q-SNAP for cobalt, dedicated to bulk, surface, and nanoparticle modeling. To demonstrate its quality, we show that it reproduces phonon dispersions and surface energies of hexagonal close-packed (hcp) and face-centered cubic (fcc) Co, as well as the relative stability of nanoparticles of various shapes at the accuracy of DFT. Dynamical properties including the $\alpha$-$\beta$ phase transition, thermal expansion, and the melting point are predicted in remarkable agreement with experimental values. Additionally, we compare the capabilities of this q-SNAP with those of one of the most widely used EAM potential for Co [24] in predicting quantities that can be critical in catalysis, such as surface energies and the vacancy formation energies on nanoparticle vertices.
The paper is organized as follows: in Section II, we outline the methods employed, beginning with a brief recapitulation of the q-SNAP formalism, followed by the approach used for fitting and testing the potential. In Section III, we demonstrate the capabilities of the new q-SNAP in terms of nanoparticle modeling. We start by analyzing the relative stability of nanoparticles of different shapes as a function of their size. Next, we compare the predictive performance of the q-SNAP and EAM potentials in determining surface energies and vacancy formation on vertices. Subsequently, we showcase the predictive capability of q-SNAP by computing the cobalt melting point through the melting of nanoparticles of increasing size, emphasizing its suitability for dynamic processes. Finally, we present our conclusions in Section IV.
@@ II Construction of the potential
II.1 SNAP formalism
An interatomic potential connects local descriptors of the atomic environment of each atom to observables such as the total energy of the entire system, forces on each atom, and stress tensors. In the case of SNAP, bispectrum descriptors are used to represent the atomic environments. In this approach, the density of neighbor atoms around a central atom $i$ at location ${\bf r}<R_{cut}$ is a sum of Dirac functions, $\delta$, located in a three-dimensional space:
$\rho_{i}({\bf r})=\delta({\bf r})+\sum_{r_{ii^{\prime}}<R_{cut}}f_{c}(r_{ii^{\prime}})w_{i^{\prime}}\delta({\bf r}-{\bf r}_{ii^{\prime}})$ (1)
where ${\bf r}_{ii^{\prime}}$ is the vector joining the two neighboring atoms $i$ and $i^{\prime}$, $w_{i^{\prime}}$ is a dimensionless coefficient used to differentiate between atoms of different types (in our case, it is always set to one since the q-SNAP is for pure cobalt) and the function $f_{c}(r)$ ensures that the contribution of each neighbor atom smoothly tends to zero at a cutoff radius $R_{cut}$.
The density is then expanded in spherical harmonics used as basis functions:
$\rho_{i}(r)=\sum_{j=0}^{\infty}\sum_{m,m^{\prime}=-j}^{j}u_{m,m^{\prime}}^{j}U_{m,m^{\prime}}^{j}(\theta,\phi,\theta_{0}),$ (2)
where $\theta_{0}$ is a third polar angle used to map the radial distance $r$ as
$\theta_{0}=\theta_{0}^{max}\frac{r}{R_{cut}}.$ (3)
The coefficients $u_{m,m^{\prime}}^{j}$ are given by the inner product of the neighbor density with the basis functions. However, these coefficients are not invariant under rotation, rendering this simple expansion in spherical harmonics not directly applicable. This issue is resolved by introducing a bispectrum of spherical harmonics, which is rotationally invariant:
$B_{j_{1},j_{2},j}=\sum_{m_{1},m^{\prime}_{1}=-j_{1}}^{j_{1}}\sum_{m_{2},m^{\prime}_{2}=-j_{2}}^{j_{2}}\sum_{m,m^{\prime}=-j}^{j}\big{(}u_{m,m^{\prime}}^{j}\big{)}$ (4)
$C_{j_{1}m_{1}j_{2}m_{2}}^{jm}\times u_{j_{1}m^{\prime}_{1},j_{2}m^{\prime}_{2}}^{jm^{\prime}}u_{m^{\prime}_{1},m_{1}}^{j_{1}}u_{m^{\prime}_{2},m_{2}}^{j_{2}},$ (5)
where $C$ are Clebsch-Gordan coefficients. These bispectrum descriptors are finally invariant under translations, rotations and permutations. $J_{max}$ is a crucial parameter that determines the number of distinct bispectrum descriptors with indices $j_{1},j_{2},j<J_{max}$. Increasing its value enhances the detailed description of the chemical environment, at the expense of computational cost. Moreover, it should be noted that excessively increasing $J_{max}$ could lead to overfitting, a topic further discussed in section II.2.2.
Once all bispectrum coefficients are determined, they are employed to compute the system’s observables. The q-SNAP energy of an atom $i$ is expressed as
$E_{\text{q-SNAP}}^{i}({\bf r}^{N})={\bf\beta}\cdot{\bf B}^{i}+\frac{1}{2}\big{(}{\bf B}^{i}\big{)}^{T}\cdot{\bf\alpha}\cdot{\bf B}^{i},$ (6)
where ${\bf B}^{i}$ are bispectrum descriptors of atom $i$ and ${\bf\alpha}$ and ${\bf\beta}$ two sets of coefficients that are optimized during the fitting process using a regression. Linear SNAP only deals with the linear part, focusing solely on optimizing the ${\bf\beta}$ coefficients. The dimension of the vectors ${\bf B}^{i}$ is related to $J_{max}$. Forces and stresses are expressed as first and second derivatives of the energy in equation 6 with respect to positions, respectively. Further details about the q-SNAP formalism can be found in Refs. 20 and 23.
II.2 Fitting process
II.2.1 Training set
Table 1: Details of the different classes used in the training set. The first column describes the class, the second one gives the number of atoms in the cells used for DFT-PBE calculations, and the three last columns correspond to the total number of energies $n_{E}$, forces $n_{F}$, and stresses $n_{S}$ used for the training/validation procedure.
@@ II.2.1 Training set
The training set is initially created by performing ab initio calculations on a set of structures that represent cobalt in all relevant configurations. Four crystallographic phases are considered: hcp, fcc, bcc and $\omega$. Unit cells of these phases are deformed in all directions up to 1% strain. Supercells and surfaces are also constructed. Various defects, such as interstitials and vacancies, are introduced into selected supercells, while adatoms, vacancies, and steps are incorporated into certain surface structures. Additionally, we generated icosahedra, truncated octahedra, cuboctahedra, decahedra, hcp, and spherical nanoparticles, each containing up to a few hundred atoms, with defects such as vacancies and adatoms also included. Subsequently, all structures are subjected to molecular dynamics simulations at temperatures up to 1200 K. Furthermore, also liquid structures are included. The EAM potential developed by Pun et. al. [24] is used to generate long trajectories with highly uncorrelated configurations. It has been found [21] that the inclusion of highly out-of-equilibrium configurations increases the robustness of the potential.
Additionally, we noticed that including small nanoclusters ranging from 2 to 30 atoms improves the potential’s robustness. Indeed, these clusters possess bispectrum descriptors that significantly differ from those of any bulk phase. In surface studies, this type of information is crucial since molecular dynamics simulations at high temperatures can easily lead to the presence of highly undercoordinated atoms, such as adatoms. A training set without these nanoclusters results in a q-SNAP that is unable to model such systems, due to an extreme increase in energy for some atoms of these special configurations.
Furthermore, we included a larger number of icosahedra compared to truncated octahedra and hcp nanoparticles, given their non-crystalline structure without any equivalent in the bulk. Otherwise, they would have been underrepresented in the training set, whereas it is often the most stable morphology for transition metal nanoparticles of a few dozen to a few hundreds atoms [25]. We chose not to increase the representation of decahedra because they are inherently unstable for cobalt [26] and fall outside the scope of our study.
@@
II.2.2 Optimization of fitting parameters
@@
Using the Machine-Learned Potential Generator (MLPG) of the MedeA materials modeling environment [31], the q-SNAP model [20, 23] was trained on the data from DFT-PBE calculations using the least squares method. To achieve the best fit, also the band limit ($2\times J_{max}$) and the radial cutoff needed to be optimized. To this end, a test set containing 10% of the training configurations was extracted from the training set, and multiple q-SNAPs with different complexities were created. For each q-SNAP, the root mean square error (RMSE) between the training and test sets was computed. The higher the complexity of the descriptors, the more the RMSE on the training set will be reduced. However, there comes a point where the error on the test set starts to increase again [34]. The optimal complexity is thus the one that minimizes the RMSE on the test set. For cobalt with our datasets, a band limit of 8 ($J_{max}=4$) and a radial cutoff of 5.0 Å are the optimal values. During the fit, weights on forces and stresses were 0.01 and 10-6 respectively. Fig. 1 shows the training and validation errors on energy and forces of the final fit. The errors are detailed in Table 2.
@@
II.3 Validation on bulk and surface properties
@@
$a$ (Å) 2.507222Ref. 35 2.4902 2.4899 2.5187
$c/a$ 1.623222Ref. 35 1.6160 1.6149 1.6103
$C_{11}$ (GPa) 293333Ref. 36, 319.50444Ref. 37, 307.1555Ref. 38 386 366 311.9
$C_{12}$ (GPa) 143333Ref. 36, 166.09444Ref. 37, 165.0555Ref. 38 151 174 146.9
$C_{13}$ (GPa) 90333Ref. 36, 102.09444Ref. 37, 102.7555Ref. 38 114 118 119.6
$C_{33}$ (GPa) 339333Ref. 36, 373.60444Ref. 37, 358.1555Ref. 38 403 425 359.4
$C_{44}$ (GPa) 78333Ref. 36, 82.41444Ref. 37, 75.5555Ref. 38 97 74 91.7
$E^{f}_{v}$ (eV) 1.4666Ref. 39, 1.38777Ref. 40 1.8222 1.49
$T_{hcp-fcc}$ (K) 690888Ref. 41, 695999Ref. 42, 700101010Ref. 43 757 717111Ref. 24
@@
$a$ (Å) 3.5447222Ref. 44, 3.568333Ref. 45 3.5103 3.5116 3.5642
$C_{11}$ (GPa) 260444Ref. 46, 225555Ref. 47, 223666Ref. 48 302 328 275.7
$C_{12}$ (GPa) 160444Ref. 46, 160555Ref. 47, 186666Ref. 48 173 201 158.9
$C_{44}$ (GPa) 110444Ref. 46, 92555Ref. 47, 110666Ref. 48 149 138 108.2
$E^{f}_{v}$ (eV) 1.34 - 1.91777Ref. 49 2.34888Ref. 50, 1.71888Ref. 50 1.7542 1.56
$T_{m}$ (K) 1768999Ref. 51, 1770101010Refs. 42 and 43 1695 +/- 15 1898111Ref. 24
We assessed the capability of the q-SNAP to accurately reproduce results obtained from periodic DFT-PBE calculations. To achieve this, numerous properties were computed and analyzed. The computational methodologies employed to calculate these properties are detailed below. DFT-PBE calculations were carried out using the parameters outlined in section II.2. For structure optimizations, we used the conjugate gradient algorithm with a convergence criterion of 0.02 eV.Å-1. The LAMMPS molecular dynamics package [52] was employed to obtain the results based on the q-SNAP. Structure minimizations were also performed using the conjugate gradient algorithm, with a convergence criterion of 10-3 eV.Å-1. During the molecular dynamics simulations, a time step of 2 fs was employed, and the Nosé-Hoover thermostat and barostat were used with dampings of 200 and 2000 fs to control temperature and pressure, respectively.
Tables 3 and 4 compare the experimental values of hcp and fcc cobalt with those obtained with q-SNAP, DFT-PBE and EAM [24]. The q-SNAP results shown in these tables reflect the accuracy of the DFT-PBE level of theory for calculations at $T=0$ K. The EAM potential was fitted to reproduce experimental data, typically measured at ambient temperature. This explains the better agreement of the EAM results for structural and elastic properties with experimental data compared with the pure DFT-PBE results. However, fitting to specific experimental data runs the risk of introducing uncontrolled errors when using such a potential to compute other properties such as the temperature of phase transitions or surface energies.
Figure 2: (Color online) Surface energies calculated with the q-SNAP and compared with those obtained by DFT-PBE and EAM potential [24] values.
@@
were $E_{slab}$ corresponds to the total energy of the slab of area of $A$ containing $n_{slab}$ atoms, and $E_{bulk_{\phi}}$ to the energy per atom of the minimized bulk cell of phase $\phi$. Slabs were constructed with a vacuum spacing of 15 Å  and a minimum thickness of 16 Å. As illustrated in Fig. 2, q-SNAP predicts surface energies for low index orientations that are extremely close to the DFT-PBE values, with a maximum difference of 47 mJ.m-2 for $\gamma_{hcp}(110)$, whereas the EAM potential leads to much larger deviations, up to 272 mJ.m-2 for the $\gamma_{fcc}(111)$. Moreover, the EAM potential finds that the hcp (001) surface is the most stable instead of the fcc (111), and predicts that $\gamma_{fcc}(011)<\gamma_{fcc}(001)<\gamma_{hcp}(110)$. The disparity observed between the EAM potential and the q-SNAP (and DFT-PBE) likely arises from the decisions made during the calibration of the EAM potential, primarily aimed at accurately reproducing bulk properties, as mentioned earlier. It should be mentioned, however, that the PBE level of theory tends to underestimate surface energies of transition metals [53].
@@
Figure 3: (Color online) Thermal expansion of hcp and fcc cobalt measured experimentally (lines) and calculated with the q-SNAP (marks). The blue solid line and squares represent the $a$ lattice parameter of hcp cobalt, whereas the dashed line and triangles refer to $c$. The red dash-dotted line and diamonds represent the lattice parameter of fcc cobalt.
@@
A further test of the present q-SNAP is the computation of the thermal expansion of hcp and fcc phases using molecular dynamics simulations in the NPT ensemble at atmospheric pressure between $T=298$ and $1200$ K. 5x5x5 and 4x4x4 supercells were employed for the hcp and fcc phases respectively. The temperature was incremented linearly in steps of 100 K with an equilibration of 50 ps for each temperature. Only the cell angles were fixed. As illustrated in Fig. 3, the q-SNAP results are in very good agreement with the experimental data from Ref. 41, showing that this interatomic potential reproduces well this property. The small underestimation of the thermal expansion at elevated temperatures may be in part due to the presence of vacancies in the experimental samples while the simulations are preformed for vacancy-free systems.
@@
Figure 4: (Color online) Phonon dispersion curves of hcp (top) and fcc (bottom) cobalt, computed using DFT-PBE (solid lines) and the q-SNAP (squares). Note that these results are obtained in the P1 space group.
@@
Phonon dispersions were computed using the phonon module as available in MedeA [31], which uses the method described in Ref. 54. The dispersions predicted by the q-SNAP are in very good agreement with those predicted by DFT-PBE, as illustrated in Fig. 4 with the highest frequencies being around 8 THz for the hcp phase, consistent with experimental data [55] while the EAM of Pun et. al. [24] predicts the highest frequencies to be nearly 11 THz.
@@
where $T$ is the temperature, $U$ the total energy of the system, $E_{vib,QC}$ the quantum contribution of the vibrational potential energy, $ZPE$ the zero point energy and $S_{vib}$ the vibrational entropy. The computations of the three last terms rely on the vibrational density of states $DOS(\nu)$, determined via the Fourier transform of the integrated velocity auto-correlation function [56, 57]. Supercells of 1000 and 1008 atoms for the hcp and fcc phases respectively are equilibrated at a given temperature for 170 ps within the NPT ensemble. Subsequently, the velocity auto-correlation function is calculated during two simulations of 25 ps in the NVE ensemble. The Helmholtz free energy of both phases is computed using equation 8, followed by polynomial fitting of the resultant curves to determine $\Delta F_{fcc-hcp}(T)$. As shown in Fig. 5, this function intersects zero at $T=757$ K, demonstrating rather good agreement with the experimental data, considering that a shift of only 1.6 meV.atom-1 would yield a transition temperature close to the experimental value of 693 K [1]. Furthermore, there might be uncertainties also in the experimental value.
The biphased cell method was used to estimate the melting temperature of fcc Co. In the NPT ensemble, following the procedure in [58], a 5x5x20 supercell was equilibrated for 100 ps at 800 K, using NPT molecular dynamics, with only the angles fixed. Subsequently, all atoms with a fractional coordinate $z$ greater than 0.5 were fixed. A 50 ps NPT molecular dynamics simulation was performed at 3000 K to rapidly transform half of the system into a liquid state. During this simulation, only the $z$-direction of the cell was allowed to relax. Then, while still keeping only the $z$-direction free, a third NPT molecular dynamics simulation was conducted at the experimental melting point of 1768 K [2]. This simulation lasted only 1 ps, which was sufficient to equilibrate the density at the presumed melting temperature without recrystallization. Once this biphased cell was constructed, NPT molecular dynamics simulations of 200 ps were performed at various temperatures. All directions were allowed to relax but the angles were fixed. The melting temperature was determined based on the final state of the final structure. All simulations were conducted at atmospheric pressure. The q-SNAP predicts a melting temperature of 1695 +/- 15 K, which is only 73 K below the experimental value.

### v2

@@ Abstract
We present the development and applications of a quadratic Spectral Neighbor Analysis Potential (q-SNAP) for ferromagnetic cobalt. Trained on Density Functional Theory calculations using the Perdew-Burke-Ernzerhof (DFT-PBE) functional, this machine-learned potential enables simulations of large systems over extended time scales across a wide range of temperatures and pressures at near DFT accuracy. It is validated by closely reproducing the phonon dispersions of hexagonal close-packed (hcp) and face-centered cubic (fcc) Co, surface energies, and the relative stability of nanoparticles of various shapes. An important feature of this novel potential is its numerical stability in long molecular dynamics simulations. This robustness is exploited to compute the heat capacity of nanoparticles containing up to 9201 atoms, showing convergence to less than 2 J.K-1.mol-1 after 100 ns. Computations of the melting temperature of nanoparticles as a function of their size revealed a convergence to the bulk limit in excellent agreement with the experimental value. Thus, the new, highly accurate machine-learned potential for Co opens exciting opportunities for further applications such as the dynamics of nanoparticles in catalytic reactions.
@@ I Introduction
Machine-learned potentials are founded on a descriptor of atomic environments and a model that links them to local energies. Descriptors are mathematical entities invariant under translation, rotation, and permutation of atoms of the same type. A diverse array of descriptors exists, such as the bispectrum (BSO(4)) [17], the Smooth Overlap of Atomic Positions (SOAP) [18], and the Atomic Cluster Expansion (ACE) [19]. Likewise, models span from simple linear regression to multi-layer neural networks [17, 20, 21, 19, 22]. Thompson et al. [20] demonstrated a linear relationship between bispectrum descriptors and system observables (energy, forces and stress), leading to the introduction of the Spectral Neighbor Analysis Potential (SNAP). Later, they showed that adding quadratic terms to the model (leading to a q-SNAP) offers improved accuracy [23].
In this article, we present the construction of a q-SNAP for cobalt, dedicated to bulk, surface, and nanoparticle modeling. To demonstrate its quality, we show that it reproduces phonon dispersions and surface energies of hexagonal close-packed (hcp) and face-centered cubic (fcc) Co, as well as the relative stability of nanoparticles of various shapes at the accuracy of DFT. Dynamical properties including the $\alpha$-$\beta$ phase transition, thermal expansion, and the melting point are predicted in remarkable agreement with experimental values. Additionally, we compare the capabilities of this q-SNAP with those of a recent EAM potential for Co [24] in predicting quantities that can be critical in catalysis, such as surface energies and the vacancy formation energies on nanoparticle vertices.
The paper is organized as follows: in Section II, we outline the approach used for fitting and testing the potential. In Section III, we validate and apply the new q-SNAP by computing properties of bulk Co, surfaces, and nanoparticles. This includes an analysis of the relative stability of nanoparticles of different shapes as a function of their size and the determination of vacancy formation on their vertices. The suitability and robustness of the new potential to model dynamic processes is demonstrated by computing the heat capacity and the dependence of the melting temperature of nanoparticles as a function of their size. The conclusions and perspectives are presented in Section IV.
@@ II Construction of the potential
II.1 Fitting process
II.1.1 Training set
Table 1: Details of the different classes used in the training set. The first column describes the class, the second one gives the number of atoms in the cells used for DFT-PBE calculations, and the three last columns correspond to the total number of energies $n_{E}$, forces $n_{F}$, and stresses $n_{S}$ used for the training/validation procedure. ”Defects” stands for bulk supercells with interstitials and vacancies.
@@ II.1.1 Training set
Figure 1: Final structure from the minimization of a cobalt cuboctahedral nanoparticle of 6525 atoms initially constructed with the lattice parameter of gold (=1.15$\times$ that of cobalt), performed with a q-SNAP trained with (right) and without (left) small nanoclusters, illustrating the need to include small nanoclusters in the training set.
While there exists a range of generic approaches for the choice of training sets such as Bayesian methods [25, 26] or entropy-based active learning [27], ultimately the choice of the training set depends on the target systems and physical phenomena such as melting, which determine the appropriate choice. Given our goal of simulating dynamic effects of bulk systems, surfaces, and especially nanoparticles, we have taken the following approach. The training set is initially created by performing ab initio calculations on a set of structures that represent cobalt in a wide range of atomic environments. Four crystallographic phases are considered: hcp, fcc, bcc and $\omega$. Even if they are not stable for cobalt, we decided to include bcc and $\omega$ phases so that the potential learns their relative stability. Unit cells of these phases are deformed in all directions up to 1% strain. Supercells and surfaces are also constructed. Various defects, such as interstitials and vacancies, are introduced into selected supercells, while adatoms, vacancies, and steps are incorporated into certain surface structures. Additionally, we generated icosahedra, truncated octahedra, cuboctahedra, decahedra, hcp, and spherical nanoparticles, each containing up to a few hundred atoms, with defects such as vacancies and adatoms also included. Furthermore, liquid structures are also included. These amorphous cells contain 60 atoms and are constructed at various densities ranking from 6.5 to 8.9 g.cm-3 using the amorphous builder of MedeA [28]. Subsequently, all these structures are subjected to molecular dynamics simulations at 400, 800 and 1200 K, using the EAM potential developed by Pun and Mishin [24]. These simulations last for 100 ps using a timestep of 2 fs. Two snapshots are included for the simulations at 400 K, whereas 3 are taken from the trajectories at 800 and 1200 K. We chose to use the EAM instead of ab initio molecular dynamics because it allows us to create long trajectories, and thus highly uncorrelated structures.
Additionally, we noticed that including small nanoclusters ranging from 2 to 30 atoms improves the potential’s robustness, as already reported in [29]. Indeed, these clusters possess bispectrum descriptors that significantly differ from those of any bulk phase. In surface studies, this type of information is crucial since molecular dynamics simulations at high temperatures can easily lead to the presence of highly undercoordinated atoms, such as adatoms. A training set without these nanoclusters results in a q-SNAP that is unable to model such systems, due to an extreme increase in energy for some atoms of these special configurations (cf. figure 1).
Finally, we included a larger number of icosahedra compared to truncated octahedra and hcp nanoparticles, given their non-crystalline structure without any equivalent in the bulk. Otherwise, they would have been underrepresented in the training set, whereas it is often the most stable morphology for transition metal nanoparticles of a few dozen to a few hundreds atoms [30]. We chose not to increase the representation of decahedra because they are inherently unstable for cobalt [31] and fall outside the scope of our study.
@@
II.1.2 Optimization of fitting parameters
@@
Using the Machine-Learned Potential Generator (MLPG) of the MedeA materials modeling environment [28], the q-SNAP model [20, 23] was trained on the data from DFT-PBE calculations using the least squares method. To achieve the best fit, also the band limit ($2\times J_{max}$) and the radial cutoff needed to be optimized. To this end, a test set containing 10% of the training configurations was extracted from the training set, and multiple q-SNAPs with different complexities were created. $J_{max}$ values of 2, 3, 4, 5 and 6 were used with radial cutoffs between 4 and 6 Å, initially by steps of 0.2 Å, then by steps of 0.1 Å around the optimized value. For each q-SNAP, the root mean square error (RMSE) between the training and test sets was computed. The higher the complexity of the descriptors, the more the RMSE on the training set will be reduced. However, there comes a point where the error on the test set starts to increase again [38]. The optimal complexity is thus the one that minimizes the RMSE on the test set. For cobalt with our datasets, a band limit of 8 ($J_{max}=4$) and a radial cutoff of 5.0 Å are the optimal values. While keeping this hyper parameters constants, weights between 0.01 and 0.1, by steps of 0.01 were tested for the forces, whereas weights of 10-5, 10-6 and 10-7 were tested for the stresses. Finally, we used weights of 1, 0.01 and 10-6 for energies, forces and stresses, respectively. Fig. 2 shows the training and validation errors on energy and forces of the final fit. The errors are detailed in Table 2.
@@
II.2 Validation on bulk and surface properties
@@
$a$ (Å) 2.507222Ref. 39 2.4902 2.4899 2.5187
$c/a$ 1.623222Ref. 39 1.6160 1.6149 1.6103
$C_{11}$ (GPa) 293333Ref. 40, 319.50444Ref. 41, 307.1555Ref. 42 386 366 311.9
$C_{12}$ (GPa) 143333Ref. 40, 166.09444Ref. 41, 165.0555Ref. 42 151 174 146.9
$C_{13}$ (GPa) 90333Ref. 40, 102.09444Ref. 41, 102.7555Ref. 42 114 118 119.6
$C_{33}$ (GPa) 339333Ref. 40, 373.60444Ref. 41, 358.1555Ref. 42 403 425 359.4
$C_{44}$ (GPa) 78333Ref. 40, 82.41444Ref. 41, 75.5555Ref. 42 97 74 91.7
$E^{f}_{v}$ (eV) 1.4666Ref. 43, 1.38777Ref. 44 1.8222 1.49
$C_{V}$ (298.15 K) (J.K-1.mol-1) 24.732888Ref. 45 22.9263999Computed within the harmonic approximation 22.9736999Computed within the harmonic approximation 22.4766999Computed within the harmonic approximation
$T_{hcp-fcc}$ (K) 690101010Ref. 46, 695111111Ref. 47, 700121212Ref. 48 757 717111Ref. 24
@@
$a$ (Å) 3.5447222Ref. 49, 3.568333Ref. 50 3.5103 3.5116 3.5642
$C_{11}$ (GPa) 260444Ref. 51, 225555Ref. 52, 223666Ref. 53 302 328 275.7
$C_{12}$ (GPa) 160444Ref. 51, 160555Ref. 52, 186666Ref. 53 173 201 158.9
$C_{44}$ (GPa) 110444Ref. 51, 92555Ref. 52, 110666Ref. 53 149 138 108.2
$E^{f}_{v}$ (eV) 1.34 - 1.91777Ref. 54 2.34888Ref. 55, 1.71888Ref. 55 1.7542 1.56
$C_{V}$ (298.15 K) (J.K-1.mol-1) 24.811999Ref. 45 23.1599101010Computed within the harmonic approximation 23.2380101010Computed within the harmonic approximation 22.6667999Ref. 45
$T_{m}$ (K) 1768111111Ref. 56, 1770121212Refs. 47 and 48 1695 +/- 15 1898111Ref. 24
We assessed the capability of the q-SNAP to accurately reproduce results obtained from periodic DFT-PBE calculations. To achieve this, numerous properties were computed and analyzed. The computational methodologies employed to calculate these properties are detailed below. DFT-PBE calculations were carried out using the parameters outlined in section II.1. For structure optimizations, we used the conjugate gradient algorithm with a convergence criterion of 0.02 eV.Å-1. The LAMMPS molecular dynamics package [57] was employed to obtain the results based on the q-SNAP. Structure minimizations were also performed using the conjugate gradient algorithm, with a convergence criterion of 10-3 eV.Å-1. During the molecular dynamics simulations, a time step of 2 fs was employed, and the Nosé-Hoover thermostat and barostat were used with dampings of 200 and 2000 fs to control temperature and pressure, respectively.
Tables 3 and 4 compare the experimental values of hcp and fcc cobalt with those obtained with q-SNAP, DFT-PBE and EAM [24]. The q-SNAP results shown in these tables reflect the accuracy of the DFT-PBE level of theory for calculations at $T=0$ K. The EAM potential was fitted to reproduce experimental data, typically measured at ambient temperature. This explains the better agreement of the EAM results for structural and elastic properties with experimental data compared with the pure DFT-PBE results. However, fitting to specific experimental data runs the risk of introducing uncontrolled errors when using such a potential to compute other properties such as the temperature of phase transitions or energies.
Figure 3: (Color online) Surface energies relative to the fcc (111) surface energy, calculated with the present q-SNAP and compared with those obtained by DFT-PBE and the EAM potential [24].
@@
were $E_{slab}$ corresponds to the total energy of the slab of area of $A$ containing $n_{slab}$ atoms, and $E_{bulk_{\phi}}$ to the energy per atom of the minimized bulk cell of phase $\phi$. Slabs were constructed with a vacuum spacing of 15 Å  and a minimum thickness of 16 Å. As illustrated in Fig. 3, q-SNAP predicts surface energies for low index orientations that are extremely close to the DFT-PBE values, with a maximum difference of 47 mJ.m-2 for $\gamma_{hcp}(110)$, whereas the EAM potential leads to much larger deviations, up to 272 mJ.m-2 for the $\gamma_{fcc}(111)$. Moreover, the EAM potential finds that the hcp (001) surface is the most stable instead of the fcc (111), and predicts that $\gamma_{fcc}(011)<\gamma_{fcc}(001)<\gamma_{hcp}(110)$. The disparity observed between the EAM potential and the q-SNAP (and DFT-PBE) likely arises from the decisions made during the calibration of the EAM potential, primarily aimed at accurately reproducing bulk properties, as mentioned above. It should be mentioned, however, that the PBE level of theory tends to underestimate surface energies of transition metals [58].
@@
Figure 4: (Color online) Thermal expansion of hcp and fcc cobalt measured experimentally (lines) and calculated with the q-SNAP (marks). The blue solid line and squares represent the $a$ lattice parameter of hcp cobalt, whereas the dashed line and triangles refer to the $c$ one. The red dash-dotted line and diamonds represent the lattice parameter of fcc cobalt.
@@
A further test of the present q-SNAP is the computation of the thermal expansion of hcp and fcc phases. To do this, we ran molecular dynamics simulations in the NPT ensemble at atmospheric pressure between $T=298$ and $1200$ K, with 5x5x5 and 4x4x4 supercells for the hcp and fcc phases respectively. The temperature was incremented linearly in steps of 100 K with an equilibration of 50 ps for each temperature. Only the cell angles were fixed. As illustrated in Fig. 4, the q-SNAP results are in very good agreement with the experimental data from Ref. 46, showing that this interatomic potential reproduces well this property. The small underestimation of the thermal expansion at elevated temperatures may be due in part to the presence of vacancies in the experimental samples while the simulations are preformed for vacancy-free systems.
@@
Figure 5: (Color online) Phonon dispersion curves of hcp (top) and fcc (bottom) cobalt, computed using DFT-PBE (solid lines), the q-SNAP (squares) and the EAM[24] (triangles). Note that these results are obtained in the P1 space group.
@@
Phonon dispersions were computed using the phonon module as available in MedeA [28], which uses the method described in Ref. 59. The dispersions predicted by the q-SNAP are in very good agreement with those predicted by DFT-PBE, as illustrated in Fig. 5, with the highest frequencies being around 8 THz for the hcp phase, consistent with experimental data [60] while the EAM of Pun et. al. [24] predicts the highest frequencies to be nearly 11 THz.
@@
where $T$ is the temperature, $U$ the internal energy of the system, $E_{vib,QC}$ the quantum contribution of the vibrational potential energy, $ZPE$ the zero point energy and $S_{vib}$ the vibrational entropy. The computations of the three last terms rely on the vibrational density of states $DOS(\nu)$, determined via the Fourier transform of the integrated velocity auto-correlation function [61, 62]. Supercells of 1000 and 1008 atoms for the hcp and fcc phases respectively are equilibrated at a given temperature for 170 ps within the NPT ensemble. Subsequently, the velocity auto-correlation function is calculated during two simulations of 25 ps in the NVE ensemble. The Helmholtz free energy of both phases is computed using equation 2, followed by polynomial fitting of the resultant curves to determine $\Delta F_{fcc-hcp}(T)$. As shown in Fig. 6, this function intersects zero at $T=757$ K, demonstrating rather good agreement with the experimental data, considering that a shift of only 1.6 meV.atom-1 would yield a transition temperature closer to the experimental value of 693 K [1]. Indeed, there might be uncertainties also in the experimental value.
Last, the biphased cell method was used to estimate the melting temperature of fcc Co. Following the procedure in [63], a 5x5x20 supercell was equilibrated for 100 ps at 800 K using NPT molecular dynamics, with only the angles fixed. Subsequently, all atoms with a fractional coordinate $z$ greater than 0.5 were fixed. A 50 ps NPT molecular dynamics simulation was performed at 3000 K to rapidly transform half of the system into a liquid state. During this simulation, only the $z$-direction of the cell was allowed to relax. Then, while still keeping only the $z$-direction free, a third NPT molecular dynamics simulation was conducted at the experimental melting point of 1768 K [2]. This simulation lasted only 1 ps, which was sufficient to equilibrate the density at the presumed melting temperature without recrystallization. Once this biphased cell was constructed, NPT molecular dynamics simulations of 200 ps were performed at various temperatures. All directions were allowed to relax but the angles were fixed. The melting temperature was determined based on the final state of the final structure. All simulations were conducted at atmospheric pressure. The q-SNAP predicts a melting temperature of 1695 +/- 15 K, which is only 73 K below the experimental value.